# Poisoning with A Pill: Circumventing Detection in Federated Learning

### v1

@@ 7.2 Limitations and Future Work
8 Conclusion
In this paper, we propose a novel attack-agnostic augmentation method to enhance existing poisoning attacks in FL by concentrating attacks into a pill (a tiny subnet). Our approach is constructed with three stages, including pill construction, pill poisoning, and pill injection. Accordingly, we first use a dynamic pill search algorithm to identify the pill blueprint. Then we poison the pill using existing FL poisoning attacks, and carefully inject the poison pill into the target model with two pill-related masks and a two-step adjustment. Our method enables existing poisoning attacks to achieve more than $2$x the error rates on average compared with their original implementations. The effectiveness of our method in exploiting and exacerbating the inherent weaknesses of current FL defenses highlights the critical need for more refined detection measures for FL.
Appendix A Additional Details of The Baseline Defenses
Krum and Multi-Krum (MKrum) [20]. Krum uses a distance score as the metric. In each round, the Krum server sums the distances between each client update $\boldsymbol{g}_{t}^{(i)}$ and its $K-m-2$ neighbors, and uses these sums as the scores for all the clients. The Krum server then selects the client’s model update with the lowest score. Multi-Krum is a variant of Krum that uses iterative Krum to pick multiple candidates for aggregation.
Coordinate-wise Median (Median) [26]. Coordinate-wise Median (Median) uses the per-parameter median values of the model updates from the clients as the aggregated global model update, which is then used to generate the next-round global model.
Trimmed Mean (Trim) [26]. Trimmed Mean (Trim) calculates per-parameter trimmed mean values of the client model updates and packs them as the global model update.
Bulyan [27]. Bulyan is a combination of Krum and Trim. It first uses the Krum-based method to select multiple candidates, and uses the per-parameter trimmed mean values of the candidate model updates as the final global model update.
FLTrust [21]. FLTrust trains a server model with a small root dataset. In each round, it computes the clipped cosine similarities between the server model update and client updates as trust scores, and then uses the trust scores as weights to aggregate all the normalized client model updates.
FLDetector (FLD) [39]. FLDetector filters out malicious clients by checking the multi-round consistency of all client updates. Malicious updates typically have lower consistency compared to benign ones.
Flame [23]. Flame utilizes HDBSCAN-based[60] dynamic clustering to filtered out malicious clients, and aggregates median-clipped benign updates with adaptive noise as the global model update.
Appendix B Additional Details of Concrete Pill Blueprints
Table VII and Table VIII illustrate the model structures of the CNN model and the simplified AlexNet, with their corresponding pill’s blueprints.
TABLE VII: Architectures of the original CNN model and the corresponding pill blueprint.
[…]
Pill Blueprint
Input $28\times 28\times 1$ $28\times 28\times 1$
Conv2d $3\times 3\times 30$ $3\times 3\times 1$
ReLU -1 -
MaxPool2d $2\times 2$ $2\times 2$
Conv2d $3\times 3\times 50$ $3\times 3\times 1$
ReLU - -
MaxPool2d $2\times 2$ $2\times 2$
Linear $1250\times 100$ $25\times 10$
ReLU - -
Linaer $100\times 10$ $10$
[…]
  • 1
  • 2
    $\times$” represents that the model does not contain this layer.
TABLE VIII: Example architectures of original AlexNet and the corresponding pill blueprint.
[…]
Pill Blueprint
Input $32\times 32\times 3$ $32\times 32\times 3$
Conv2d $11\times 11\times 48$ $11\times 11\times 1$
ReLU - -
MaxPool2d $3\times 3$ $3\times 3$
Conv2d $3\times 3\times 96$ $3\times 3\times 1$
ReLU - -
MaxPool2d $3\times 3$ $3\times 3$
Conv2d $3\times 3\times 192$ $3\times 3\times 1$
ReLU - -
Conv2d $3\times 3\times 192$ $3\times 3\times 1$
ReLU - -
Conv2d $3\times 3\times 128$ $3\times 3\times 1$
ReLU - -
MaxPool2d $3\times 3$ $3\times 3$
Linear $4608\times 1024$ $36\times 1$
ReLU - -
Linear $1024\times 512$ $1\times 10$
ReLU - -
Linear $512\times 10$ $10$
Softmax - $\times$
Appendix C Additional Details of The Dynamic Patterns in Our Method
Considering advantages and disadvantages of both the one-time searching strategy and the repeated searching strategy, we design a more flexible searching strategy, termed as “adaptive searching strategy”. In adaptive searching strategy, our method searches the new pill only when the pill is not successfully injected into the global model in the last round. The condition:
[…]
should be satisfied to trigger the new subnetwork searching on malicious client $i$, where $C_{search}$ is set as $0.94$ in the experiments. The adaptive searching strategy is a mroe moderate version of repeated searching.
Since the three searching strategies have their unique advantages, we investigate different combinations of them in the experiments. We further divide the neural network into Feature Extractor (FE) and Classifier (CLS). Refer to the CNN model we used, the convolutional layers are regarded as FE, and the linear layers are regarded as CLS. We use different searching strategies in FE and CLS, respectively. In all the nine combinations, we test and keep six of them, noted as Pattern1 to Pattern6, shown in Table II in § 5.2. Such six patterns construct the entire dynamic pattern set used in our method.
TABLE IX: Error rates under cross-silo setting using “approximate max pill search” (20% malicious clients) on MNIST dataset.
@@ Appendix C Additional Details of The Dynamic Patterns in Our Method
Sign-flipping Attack 0.934 0.059 0.038 0.055 0.055 0.036 0.025 0.886 0.073 0.041 0.052 0.059 0.041 0.026
+ Poison Pill 0.353 0.093 0.454 0.283 0.268 0.173 0.588 0.431 0.059 0.605 0.349 0.333 0.217 0.713
+ Poison Pill 0.416 0.109 0.469 0.252 0.247 0.117 0.026 0.581 0.065 0.672 0.358 0.324 0.092 0.051
+ Poison Pill 0.326 0.082 0.585 0.266 0.272 0.169 0.632 0.528 0.062 0.556 0.350 0.321 0.210 0.746
+ Poison Pill 0.402 0.106 0.518 0.273 0.262 0.218 0.766 0.534 0.077 0.707 0.369 0.318 0.194 0.861
TABLE X: Error rates under cross-device setting using “approximate max pill search” (20% malicious clients) on Fashion-MNIST dataset in both IID and non-IID data distribution.
Data Distribution IID Non-IID
$\pm$0.004 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.010 $\pm$0.003 $\pm$0.002 $\pm$0.008 $\pm$0.006 $\pm$0.002 $\pm$0.009 $\pm$0.005
Sign-flipping Attack 0.940 0.116 0.110 0.128 0.165 0.121 0.905 0.124 0.118 0.136 0.184 0.134
$\pm$0.026 $\pm$0.003 $\pm$0.003 $\pm$0.005 $\pm$0.007 $\pm$0.004 $\pm$0.031 $\pm$0.004 $\pm$0.002 $\pm$0.003 $\pm$0.007 $\pm$0.006
0.591 0.117 0.749 0.357 0.589 0.225 0.573 0.125 0.665 0.379 0.662 0.277
+ Poison Pill $\pm$0.177 $\pm$0.004 $\pm$0.076 $\pm$0.057 $\pm$0.048 $\pm$0.026 $\pm$0.140 $\pm$0.004 $\pm$0.111 $\pm$0.018 $\pm$0.131 $\pm$0.012
$\pm$0.018 $\pm$0.004 $\pm$0.010 $\pm$0.002 $\pm$0.014 $\pm$0.004 $\pm$0.048 $\pm$0.002 $\pm$0.025 $\pm$0.016 $\pm$0.016 $\pm$0.011
0.620 0.492 0.620 0.228 0.424 0.232 0.654 0.533 0.679 0.324 0.483 0.226
+ Poison Pill $\pm$0.051 $\pm$0.023 $\pm$0.025 $\pm$0.025 $\pm$0.042 $\pm$0.035 $\pm$0.037 $\pm$0.041 $\pm$0.049 $\pm$0.098 $\pm$0.098 $\pm$0.025
$\pm$0.002 $\pm$0.004 $\pm$0.010 $\pm$0.005 $\pm$0.009 $\pm$0.003 $\pm$0.005 $\pm$0.004 $\pm$0.031 $\pm$0.011 $\pm$0.014 $\pm$0.006
0.681 0.138 0.740 0.362 0.572 0.258 0.604 0.141 0.750 0.372 0.649 0.277
+ Poison Pill $\pm$0.057 $\pm$0.015 $\pm$0.092 $\pm$0.073 $\pm$0.167 $\pm$0.018 $\pm$0.125 $\pm$0.009 $\pm$0.081 $\pm$0.035 $\pm$0.184 $\pm$0.013
$\pm$0.005 $\pm$0.002 $\pm$0.036 $\pm$0.012 $\pm$0.011 $\pm$0.006 $\pm$0.014 $\pm$0.024 $\pm$0.112 $\pm$0.072 $\pm$0.013 $\pm$0.007
0.651 0.244 0.718 0.312 0.503 0.249 0.670 0.229 0.621 0.349 0.581 0.386
+ Poison Pill $\pm$0.082 $\pm$0.104 $\pm$0.059 $\pm$0.026 $\pm$0.060 $\pm$0.014 $\pm$0.123 $\pm$0.098 $\pm$0.030 $\pm$0.047 $\pm$0.161 $\pm$0.141
TABLE XI: Error rates under cross-device setting using “approximate max pill search” (10% malicious clients) on Fashion-MNIST dataset in both IID and non-IID data distribution.
Data Distribution IID Non-IID
$\pm$0.004 $\pm$0.004 $\pm$0.005 $\pm$0.003 $\pm$0.008 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.003 $\pm$0.002 $\pm$0.004 $\pm$0.002
Sign-flipping Attack 0.929 0.111 0.108 0.111 0.153 0.117 0.902 0.118 0.115 0.120 0.175 0.134
$\pm$0.026 $\pm$0.004 $\pm$0.003 $\pm$0.002 $\pm$0.025 $\pm$0.004 $\pm$0.034 $\pm$0.001 $\pm$0.004 $\pm$0.003 $\pm$0.008 $\pm$0.008
0.195 0.114 0.170 0.138 0.347 0.137 0.330 0.124 0.165 0.148 0.483 0.161
+ Poison Pill $\pm$0.032 $\pm$0.003 $\pm$0.071 $\pm$0.005 $\pm$0.059 $\pm$0.004 $\pm$0.135 $\pm$0.006 $\pm$0.016 $\pm$0.007 $\pm$0.225 $\pm$0.009
$\pm$0.003 $\pm$0.004 $\pm$0.002 $\pm$0.005 $\pm$0.021 $\pm$0.004 $\pm$0.003 $\pm$0.004 $\pm$0.004 $\pm$0.003 $\pm$0.011 $\pm$0.003
0.369 0.138 0.212 0.128 0.310 0.140 0.589 0.154 0.300 0.139 0.351 0.156
+ Poison Pill $\pm$0.147 $\pm$0.015 $\pm$0.066 $\pm$0.005 $\pm$0.038 $\pm$0.014 $\pm$0.046 $\pm$0.017 $\pm$0.100 $\pm$0.006 $\pm$0.056 $\pm$0.015
$\pm$0.003 $\pm$0.001 $\pm$0.003 $\pm$0.004 $\pm$0.004 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.004 $\pm$0.008 $\pm$0.001
0.164 0.117 0.157 0.143 0.371 0.142 0.229 0.126 0.249 0.146 0.374 0.157
+ Poison Pill $\pm$0.038 $\pm$0.003 $\pm$0.044 $\pm$0.010 $\pm$0.034 $\pm$0.005 $\pm$0.069 $\pm$0.002 $\pm$0.167 $\pm$0.004 $\pm$0.023 $\pm$0.005
$\pm$0.002 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.006 $\pm$0.003 $\pm$0.002 $\pm$0.001 $\pm$0.004 $\pm$0.004 $\pm$0.008 $\pm$0.001
0.351 0.124 0.299 0.135 0.343 0.146 0.342 0.138 0.292 0.148 0.417 0.166
+ Poison Pill $\pm$0.204 $\pm$0.019 $\pm$0.110 $\pm$0.004 $\pm$0.070 $\pm$0.019 $\pm$0.076 $\pm$0.018 $\pm$0.087 $\pm$0.009 $\pm$0.050 $\pm$0.010
Appendix D Additional Stealthiness Analysis
As shown in Figure 5, in addition to the minimal difference between the benign updates and the augmented malicious updates, our method also achieves two more improvements. First, our method causes the global model to degrade earlier compared to the original attacks, further demonstrating the effectiveness of our augmentation. Second, our method significantly increases the discrepancy between benign client updates as the communication rounds increase. While original attacks can bypass detection in some cases, the discrepancy between benign client updates remains steady, illustrating the lower impact of malicious clients. In contrast, our method consistently increases the discrepancy among benign clients, highlighting its penetrating effectiveness in its influence on benign clients’ local training.
Appendix E Additonal Details on MNIST and CIFAR-10 Dataset
The detailed results on MNIST and CIFAR-10 dataset are presented in Table IX (MNIST dataset) and Table V (CIFAR-10 dataset), respectively.
On CIFAR-10 dataset, our method help existing FL poisoning attacks outperform their original versions in $71$ of the $72$ scenarios, with an average error rate increase of over $0.288$. Specifically, our method facilitates at least a $0.212$ increase in error rates against FLTrust, outperforming the results in the same settings on the Fashion-MNIST dataset.
Appendix F Additional Results in Cross-device FL System
After evaluating our method in the 50-client cross-silo FL system, we further test it in the 50-client cross-device FL system. Table X presents the error rates under the cross-device FL setting using the “approximate max pill search” algorithm on both IID and non-IID data. We report the highest error rates among the results of six dynamic patterns, with the malicious client proportion set to $20\%$. Since FLD is not typically designed for cross-device systems, we do not test it in this setting.
Results on IID Data. The highest error rate improvement with our method achieves $0.639$, and the average error rate increase with our method reaches $0.279$. With our method, existing model poisoning attacks outperform their original versions in $22$ out of the $24$ cases. The highest error rate improvement for the sign-flipping attack is $0.639$, with an average error rate increase of $0.279$. For the Trim attack and Krum attack, the highest error rate increases are $0.469$ and $0.568$, with average error rate increases of $0.264$ and $0.302$, respectively. For the Min-Max attack, the highest error rate increase reaches $0.505$, with an average increase of $0.272$. These improvements are consistent with the error rates observed under the cross-silo FL setting using the "approximate max pill search" algorithm on IID data.
Results on non-IID Data. As for the results on non-IID data, the highest error rate improvement with our method achieves $0.546$, and the average error rate increase with our method reaches $0.273$. By using our method, existing model poisoning attacks outperform their original versions in 21 out of the 24 cases. The highest error rate improvement for the sign-flipping attack is $0.547$, with an average error rate increase of $0.282$. For the Trim attack and Krum attack, the highest error rate rises are $0.451$ and $0.546$, with average error rate rises of $0.312$ and $0.278$, respectively. For the Min-Max attack, the highest error rate increase reaches $0.479$, with an average increase of $0.201$. These improvements are also aligned with the error rates observed under the cross-silo FL setting using the max subnetwork searching algorithm on non-IID data.
Appendix G Additional Results with Fewer Malicious Clients
We also test the error rate improvement of our method in both the IID and non-IID cross-device FL systems, with only $10\%$ malicious clients. The experimental results are shown in Table XI.

### v2

@@ Appendix I Additional Augmentation Performance Analysis
Appendix Q Additional Ablation Study on Pill Adjustment
To illustrate the necessity of both the SimAdjust() and DistAdjust() used in our method, we conduct a detailed ablation study, providing the error rates of the Sign-flipping Attack and the Trim Attack with different settings of pill adjustments on the IID and non-IID Fashion-MNIST dataset within a 50-client FL system in Table 12.
The results show that using both SimAdjust() and DistAdjust() together achieves the highest error rate in multiple cases. For Sign-flipping Attack, the combined adjustment method outperforms others against multi-Krum (both IID and non-IID), FLDetector (both IID and non-IID), and Bulyan (non-IID). Similarly, for Trim Attack, the combined approach yields the highest error rates in 8 out of 14 defenses under both IID and Non-IID settings.
In summary, the results confirm that SimAdjust() and DistAdjust() complement each other, improving attack effectiveness while maintaining stealthiness when used together.
Appendix R Additional Results with Label-flipping Attack
In addition to the untargeted model poisoning attacks discussed in the main text, we evaluate our method using a data-poisoning-based targeted attack: the label-flipping attack. Label-flipping is a straightforward yet effective targeted attack in federated learning (FL) and is also among the least stealthy data-poisoning-based attacks. To make the evaluation more challenging, we configured the attacker to flip all the labels of the training data on malicious clients, making the label-flipping attack even less stealthy. The results are shown in Table 13, demonstrating that our method enhances the label-flipping attack to bypass six additional defenses compared to its original version. This illustrates the compatibility of our method for data-poisoning-based and targeted attacks.
Appendix S Additional Results with More Complex Models
To further demonstrate the effectiveness of our method on more complex model architectures, we test our method using VGG-11 net on the IID CIFAR-10 dataset, shown in Table 14. The results demonstrate that our method consistently enhances the performance of Sign-flipping Attack, Trim Attack, and Krum Attack, outperforming their original versions in 41 out of 54 cases. These findings illustrate the effectiveness of our method with more complex model architectures.
Appendix T Additional Results with Larger FL Systems
To further demonstrate the effectiveness of our method in larger FL systems, we extend our experiments on the Fashion-MNSIT dataset with 100 clients, shown in Table 15. The results show a similar trend as observed in the 50-client system. Our method enables baseline attacks to successfully bypass four additional defenses, causing over 50% additional error rates in the global model. These findings further validate the effectiveness and generality of our approach in larger systems, when a single malicious client has fewer data samples.
Appendix U Comparison with Existing Attack Enhancement Method
Considering several prior studies (bagdasaryan2020backdoor; bhagoji2019analyzing; zhang2022neurotoxin) enhancing backdoor attacks, we also adapt one recent one - Neurotoxin (zhang2022neurotoxin) - to our untargeted attacks evaluation setting. Table 16 illustrates the results on the IID Fashion-MNIST dataset within a 50-client FL system using Trim Attack. The results demonstrate that our method outperforms Neurotoxin in 24 out of 28 cases. This highlights that directly transferring existing methods designed for backdoor attacks may not yield consistent effectiveness when applied to untargeted attack scenarios. The results further validate the robustness of our approach.
@@
Sign-Flipping 0.935 0.123 0.098 0.101 0.106 0.101 0.103 0.903 0.114 0.103 0.107 0.124 0.111 0.116
[…]
Trim Attack 0.102 0.112 0.101 0.106 0.113 0.103 0.095 0.118 0.113 0.118 0.113 0.136 0.114 0.108
[…]
Table 17: Performance when the number of malicious clients is gradually decreasing.
Appendix V Results with Decreasing Number of Malicious Clients
To further demonstrate the effectiveness of our method in a more practical setting, we evaluate its performance as the number of malicious clients in the FL system gradually decreases. Specifically, we used the Fashion-MNIST dataset within a 50-client FL system. Initially, 20% of clients are malicious, and for every T/4 rounds (where T is the total number of FL communication rounds), the proportion of malicious clients reduces by 5%. Here is a detailed breakdown of this setup:
  • $0\rightarrow\frac{T}{4}$: 20% clients in the FL system are malicious.
  • $\frac{T}{4}\rightarrow\frac{T}{2}$: 15% clients in the FL system are malicious.
  • $\frac{T}{2}\rightarrow\frac{3T}{4}$: 10% clients in the FL system are malicious.
  • $\frac{3T}{4}\rightarrow T$: 5% clients in the FL system are malicious.
The results are presented in Table 17, demonstrating that our method significantly enhances the error rates achieved by the original Trim Attack and Sign-flipping attack in 23 out of 28 cases, with an average error rate increase of over 50%. These findings illustrate the robustness and effectiveness of our method in a more practical scenario.